one angle of a triangle is twice as large as another and 25 degree more than the third angle. find the angles.
Answer by sachi(548)

3rd angle=x
1st angle=x+25
2nd angle=[x+25]/2
now x+x+25+[x+25]/2=180
or 5x+50+25=180*2=360
or 5x=360-75=285
or x=57
so the angles are 57,41,82
ans